# On a conjecture of Peter Neumann on fixed points in permutation groups

### v1

@@ 6.7. Almost simple subgroups in class $\mathcal{U}$}
Now $r$ must divide the order of a maximal torus of $G$, which implies that $q_{0}\leq q^{8}$. From (6.4) and inspection of semisimple element centralizers in $E_{8}(q)$ (see [19]), it follows that $C_{\bar{G}}(g)=E_{7}T_{1}$, $D_{7}T_{1}$, $E_{6}A_{1}T_{1}$ or $E_{6}T_{2}$, where $T_{i}$ denotes a torus of rank $i$. In the $E_{6}T_{2}$ case we have $g\in T_{2}$ and so $q_{0}\leq q^{2}$; but then (6.4) does not hold. Hence
@@
where $\tau=\frac{1}{2}(1+\sqrt{5})$. From [35], we see that the 3-modular irreducibles for $Alt_{6}$ are $V_{1}$, $V_{3}$, $V_{3}^{*}$, $V_{4}$ and $V_{9}$, where $V_{i}$ has dimension $i$. Let their multiplicities as composition factors of $L(E_{8})\downarrow M_{0}$ be $a,b,c,d,e$ respectively. Then the values of $\chi(1)$, $\chi(x)$ and $\chi(y)$ give the equations
@@ 7.1. Some representation theory
Let $G$ be a finite group and let $W$ be an indecomposable $KG$-module. Let $P$ be a cyclic Sylow $p$-subgroup of $G$, let $S$ be an abelian subgroup of $G$ containing $P$, and assume that
@@
By [6, Corollary 3.6.10], every $KG$-module is projective relative to $P$, i.e., it is a direct summand of some module induced from $P$ (see [6, Proposition 3.6.4]). Moreover, assumption ($\star$) implies that
@@ 7.2. Some optimization
The following few lemmas will be used to bound the dimension $d$ of the centralizer of certain semisimple elements $g\in\mathrm{GL}_{n}(K)=\mathrm{GL}(V)$. For example, assume that $\dim(C_{V}(g))\leq n/3$ and that the dimensions of the nontrivial eigenspaces of $g$ occur with multiplicity at least $4$ (we will use Lemma 7.1 in order to produce such elements). Then Lemma 7.6, below, applied with $A=n/3$ and $B=4$, asserts that $d\leq n^{2}/4$; equality is attained when $C_{V}(g)=0$ and $g$ has four nontrivial eigenspaces of dimension $n/4$. As another example, assuming that $\dim(C_{V}(g))\leq n/2$ and that the dimensions of the nontrivial eigenspaces of $g$ occur with multiplicity at least $4$, we have $d\leq 5n^{2}/16$, with equality attained when $C_{V}(g)$ has dimension $n/2$ and there are four nontrivial eigenspaces of dimension $n/8$.
@@
Let $B\geq 2,C\geq 0$, and $n$ be integers, with $B,C\leq n$, and let $a_{1}\geq\cdots\geq a_{n}$ be a partition of $n$. Assume that there exists a subset $I$ of $\{1,\ldots,n\}$ such that $\sum_{i\in I}a_{i}\leq C$, and $a_{j}\geq B$ for every $j\not\in I$. Then
@@
The case $C=n$ is trivial, so assume $C<n$. Fix any $C^{\prime}\leq C$, and assume $\sum_{i\in I}a_{i}=C^{\prime}$. Assume that $a_{i}>1$ for some $i\in I$; then consider the partition $b_{1}\geq\cdots\geq b_{n}$ obtained by replacing $a_{i}$ by a part of length $a_{i}-1$, and adding one part of length $1$. In the transpose partition, this means that $b^{\prime}_{1}=a^{\prime}_{1}+1$, and $b^{\prime}_{j}=a^{\prime}_{j}-1$ for some $j>1$. Since
@@
Assume first $C^{\prime}=C$. The condition $a_{j}\geq B$ for every $j\not\in I$ is equivalent to $a^{\prime}_{1}-C=a^{\prime}_{2}=\cdots=a^{\prime}_{B}$. (Note that $I$ is a proper subset of $\{1,\ldots,n\}$ since $C<n$.) Assume that there exists $j>B$ with $0<a^{\prime}_{j}<a^{\prime}_{1}-C$; without loss, $j=B+1$. Note that
@@ 7.3. Centralizer bounds
    Assume $Y=\mathrm{Sp}_{n}(K)$. If $g$ is unipotent and $p$ is odd, then $d^{\prime}\leq d/2+R/2$ where $R$ is the number of Jordan blocks of $g$. If $g$ is semisimple, without eigenvalue 1-1 if $p$ is odd, then $d^{\prime}=d/2+m/2$ where $m$ is the dimension of the $1$-eigenspace of $g$.
@@
    Assume $Y=\mathrm{O}_{n}(K)$. If $g$ is unipotent and $p$ is odd, then $d^{\prime}\leq d/2$. If $g$ is semisimple, without eigenvalue 1-1 if $p$ is odd, then $d^{\prime}=d/2-m/2$ where $m$ is the dimension of the $1$-eigenspace of $g$.
@@
    Assume $g$ is semisimple, and let $E$ (resp. $E_{2}$) be the number of distinct irreducible factors (resp. irreducible factors of degree at least $2$) of the characteristic polynomial of $g$. If $G=\mathrm{Sp}_{n}(q)$ then $|C_{G}(g)|<2^{E_{2}}q^{d^{\prime}}\leq q^{d^{\prime}+E_{2}}$; if $G=\mathrm{O}^{\varepsilon}_{n}(q)$ then $|C_{G}(g)|<2^{E_{2}+2}q^{d^{\prime}}\leq q^{d^{\prime}+E_{2}+2}$; and if $G=\mathrm{GU}_{n}(q)$ then $|C_{G}(g)|<2^{E}q^{d^{\prime}}\leq q^{d^{\prime}+E}$.
@@
Let now $g_{f}$ be the restriction of $g$ to $V_{f}$. Assume $f$ has degree $m$ and $V_{f}$ has dimension $k$. Note that $g_{f}$ acts $\mathbb{F}_{q^{m}}$-linearly; let $u_{f}$ be the unipotent part of $g_{f}$ as an element of $\mathrm{GL}_{k/m}(q^{m})$. Write $S=S_{f,f^{*}}$, $g^{\prime}=g_{f,f^{*}}$, $u=u_{f}$, for ease of notation. Then $C_{S}(g^{\prime})\cong C_{L}(u)$, where $L=\mathrm{GU}_{k/m}(q^{m/2})$ if $f=f^{*}$ and $\deg(f)>1$; $L=\mathrm{GL}_{k/m}(q^{m})$ if $f\neq f^{*}$; $L=\mathrm{Sp}_{k}(q)$ if $f=f^{*}$ and $\deg(f)=1$ (i.e., $f=X\pm 1$). Assume next $g$ is as in (ii). Then $g$ is semisimple, and so $u=1$ and $C_{L}(u)=L$. In particular, we see that the bound follows from Lemma 7.9, noting that the number of $f$ with $f=f^{*}$ and $f\neq X+1$ and $V_{f}\neq 0$ is at most $E_{2}$ (since each such $f$ has degree at least $2$). Assume now $g$ is as in (iii), so $u$ may be nontrivial. The bound follows from (i), noting that the number of rational canonical blocks of $g^{\prime}$ is at least the number of Jordan blocks of $u$. (More precisely, if $f=f^{*}$ then it is equal to the number of Jordan blocks of $u$, and otherwise it is twice that number.)
@@
Let us then focus on Table 7.1. We denote by $\Phi_{m}(X)$ the $m$-th cyclotomic polynomial, and by $\Phi^{*}_{m}(r)$ the product of all primitive prime divisors (ppds) of $r^{m}-1$. Recall that by Zsigmondy’s theorem, $\Phi^{*}_{m}(r)>1$ unless $m=2$ and $r+1$ is a $2$-power, or $(m,r)=(6,2)$. (The notation $\Phi^{*}_{m}(r)$, which is borrowed from [27], is not accidental, in the sense that $\Phi^{*}_{m}(r)$ always divides the $m$-th cyclotomic polynomials $\Phi_{m}$ evaluated at $r$.) We also denote $\Phi^{\prime}_{8}=r+\sqrt{2r}+1$, $\Phi^{\prime}_{12}=r+\sqrt{3r}+1$, $\Phi^{\prime}_{24}=r^{2}+\sqrt{2r^{3}}+r+\sqrt{2r}+1$.
@@
Note that in some cases the order of $x_{i}$ is $1$. For $x_{1}$, this happens when $S$ is $\mathrm{PSL}_{6}(2)$ or $\mathrm{P}\Omega^{+}_{8}(2)$. For $x_{2}$, this happens when $S$ is $\mathrm{PSL}_{7}(2)$, $\mathrm{PSp}_{4}(r)$ (and $r+1$ is a $2$-power), $\mathrm{PSp}_{8}(2)$, $\mathrm{P}\Omega^{+}_{12}(2)$, or $\mathrm{P}\Omega^{-}_{8}(2)$. In particular, in all these cases we regard $x_{i}$ as being not defined.
@@
(i) The last statement (“and so…”) follows from the first. The first statement follows from inspection of the Schur multiplier $M$ of $S$ and, in groups of Lie type, from the fact that each prime divisor of $\Phi^{*}_{c}(r)$ is congruent to $1$ mod $c$. Let us handle for example $S=\mathrm{PSL}_{m}(r)$. Each prime divisor of $M$ divides $m$, and each prime divisor of $|x_{1}|$ is at least $m+1$, so $(|x_{1}|,M)=1$. As for $x_{2}$, the only possible exception occurs when $m$ is a prime divisor of $\Phi^{*}_{m-1}(r)$. But then by definition of ppd, we have that $m$ does not divide r1r-1 and so does not divide $M$.
@@
(iii) Assume first $S$ is classical, let $\tilde{S}$ be the universal covering group of $S$ and let $\tilde{x}_{i}$ be a lift of $x_{i}$ as in (i). We readily see that $\tilde{x}_{i}$ has distinct eigenvalues on the natural module (in which case $x_{i}$ is regular) except possibly in the following cases: $i=2$ and $S=\mathrm{PSU}_{m}(r),\mathrm{PSp}_{2m}(r),\mathrm{P}\Omega^{\pm}_{2m}(r),\mathrm{P}\Omega_{2m+1}(r)$; or $i=1$ and $S=\mathrm{P}\Omega^{+}_{2m}(r)$. In the case $i=2$ and $S=\mathrm{P}\Omega^{-}_{2m}(r)$, or $i=1$ and $S=\mathrm{P}\Omega^{+}_{2m}(r)$, we have that $\tilde{x}_{i}$ centralizes a $2$-space and acts irreducibly on a complement, so $x_{i}$ is regular. In the case $i=2$ and $S=\mathrm{PSp}_{2m}(r)$, or $S=\mathrm{PSU}_{m}(r)$ with $m\equiv 3\pmod{4}$, $\tilde{x}_{i}$ centralizes a $2$-space and so $x_{i}$ is not regular. Let us go through the remaining cases. If $S=\mathrm{PSU}_{m}(r)$ with $m$ even or $m\equiv 1\pmod{4}$, a straightforward calculation shows that $\tilde{x}_{2}$ has distinct eigenvalues if and only if $m\equiv 0,1\pmod{4}$ (that is, if and only if a totally singular irreducible $\mathbb{F}_{r}\langle\tilde{x}_{2}\rangle$-submodule has even dimension). On the other hand, if $S=\mathrm{P}\Omega^{+}_{2m}(r),\mathrm{P}\Omega_{2m+1}(r)$ then $\tilde{x}_{2}$ has distinct eigenvalues if and only if $m$ is odd. This concludes the proof if $S$ is classical.
@@
(vii) If $S$ is sporadic we use [12], and if $S$ is of Lie type this is a straightforward check. We only note that the parameter $\varepsilon$ for $S=G_{2}(r)$ ensures that $|x_{1}|$ is not divisible by $3$. ∎
@@
Let us recall that if $x_{i}$ is regular, then the value $N_{L}(\langle y_{i}\rangle)/C_{L}(y_{i})$ can be read off from the Weyl group, as follows. Assume for convenience that $L=X^{F}$ is of simply connected type, so by Lemma 7.12(iv) we have $N_{L}(\langle y_{i}\rangle)/C_{L}(y_{i})=N_{X^{F}}(T^{F})/T^{F}$. Write $T=R_{w}$ where $R$ is a fixed $F$-stable maximal torus and $w\in W:=N_{X}(R)/R$ (see [63, Section 25.1] for the notation used). Since in our case $N_{X^{F}}(T^{F})=N_{X^{F}}(T)$, we have that $|N_{X^{F}}(T^{F})/T^{F}|=|C_{W}(Fw)|$, where we view $Fw$ as an element of the coset $FW$ of $W\rtimes\langle F\rangle$; see for example [63, Proposition 25.3].

### v2

@@ 6.7. Almost simple subgroups in class $\mathcal{U}$}
Now $r$ must divide the order of a maximal torus of $G$, which implies that $q_{0}\leq q^{8}$. From (6.4) and inspection of semisimple element centralizers in $E_{8}(q)$ (see [19]), it follows that $C_{\bar{G}}(g)=E_{7}T_{1}$, $D_{7}T_{1}$, $E_{6}A_{1}T_{1}$ or $E_{6}T_{2}$, where $T_{i}$ denotes a torus of rank ii. In the $E_{6}T_{2}$ case we have $g\in T_{2}$ and so $q_{0}\leq q^{2}$; but then (6.4) does not hold. Hence
@@
where $\tau=\frac{1}{2}(1+\sqrt{5})$. From [35], we see that the 3-modular irreducibles for $Alt_{6}$ are $V_{1}$, $V_{3}$, $V_{3}^{*}$, $V_{4}$ and $V_{9}$, where $V_{i}$ has dimension ii. Let their multiplicities as composition factors of $L(E_{8})\downarrow M_{0}$ be $a,b,c,d,e$ respectively. Then the values of $\chi(1)$, $\chi(x)$ and $\chi(y)$ give the equations
@@ 7.1. Some representation theory
Let $G$ be a finite group and let $W$ be an indecomposable KGKG-module. Let $P$ be a cyclic Sylow $p$-subgroup of $G$, let $S$ be an abelian subgroup of $G$ containing $P$, and assume that
@@
By [6, Corollary 3.6.10], every KGKG-module is projective relative to $P$, i.e., it is a direct summand of some module induced from $P$ (see [6, Proposition 3.6.4]). Moreover, assumption ($\star$) implies that
@@ 7.2. Some optimization
The following few lemmas will be used to bound the dimension $d$ of the centralizer of certain semisimple elements $g\in\mathrm{GL}_{n}(K)=\mathrm{GL}(V)$. For example, assume that $\dim(C_{V}(g))\leq n/3$ and that the dimensions of the nontrivial eigenspaces of $g$ occur with multiplicity at least 44 (we will use Lemma 7.1 in order to produce such elements). Then Lemma 7.6, below, applied with $A=n/3$ and $B=4$, asserts that $d\leq n^{2}/4$; equality is attained when $C_{V}(g)=0$ and $g$ has four nontrivial eigenspaces of dimension $n/4$. As another example, assuming that $\dim(C_{V}(g))\leq n/2$ and that the dimensions of the nontrivial eigenspaces of $g$ occur with multiplicity at least 44, we have $d\leq 5n^{2}/16$, with equality attained when $C_{V}(g)$ has dimension $n/2$ and there are four nontrivial eigenspaces of dimension $n/8$.
@@
Let $B\geq 2,C\geq 0$, and $n$ be integers, with $B,C\leq n$, and let $a_{1}\geq\cdots\geq a_{n}$ be a partition of $n$. Assume that there exists a subset II of $\{1,\ldots,n\}$ such that $\sum_{i\in I}a_{i}\leq C$, and $a_{j}\geq B$ for every $j\not\in I$. Then
@@
The case $C=n$ is trivial, so assume $C<n$. Fix any $C^{\prime}\leq C$, and assume $\sum_{i\in I}a_{i}=C^{\prime}$. Assume that $a_{i}>1$ for some $i\in I$; then consider the partition $b_{1}\geq\cdots\geq b_{n}$ obtained by replacing $a_{i}$ by a part of length $a_{i}-1$, and adding one part of length 11. In the transpose partition, this means that $b^{\prime}_{1}=a^{\prime}_{1}+1$, and $b^{\prime}_{j}=a^{\prime}_{j}-1$ for some $j>1$. Since
@@
Assume first $C^{\prime}=C$. The condition $a_{j}\geq B$ for every $j\not\in I$ is equivalent to $a^{\prime}_{1}-C=a^{\prime}_{2}=\cdots=a^{\prime}_{B}$. (Note that II is a proper subset of $\{1,\ldots,n\}$ since $C<n$.) Assume that there exists $j>B$ with $0<a^{\prime}_{j}<a^{\prime}_{1}-C$; without loss, $j=B+1$. Note that
@@ 7.3. Centralizer bounds
    Assume $Y=\mathrm{Sp}_{n}(K)$. If $g$ is unipotent and $p$ is odd, then $d^{\prime}\leq d/2+R/2$ where $R$ is the number of Jordan blocks of $g$. If $g$ is semisimple, without eigenvalue 1-1 if $p$ is odd, then $d^{\prime}=d/2+m/2$ where $m$ is the dimension of the 11-eigenspace of $g$.
@@
    Assume $Y=\mathrm{O}_{n}(K)$. If $g$ is unipotent and $p$ is odd, then $d^{\prime}\leq d/2$. If $g$ is semisimple, without eigenvalue 1-1 if $p$ is odd, then $d^{\prime}=d/2-m/2$ where $m$ is the dimension of the 11-eigenspace of $g$.
@@
    Assume $g$ is semisimple, and let $E$ (resp. $E_{2}$) be the number of distinct irreducible factors (resp. irreducible factors of degree at least 22) of the characteristic polynomial of $g$. If $G=\mathrm{Sp}_{n}(q)$ then $|C_{G}(g)|<2^{E_{2}}q^{d^{\prime}}\leq q^{d^{\prime}+E_{2}}$; if $G=\mathrm{O}^{\varepsilon}_{n}(q)$ then $|C_{G}(g)|<2^{E_{2}+2}q^{d^{\prime}}\leq q^{d^{\prime}+E_{2}+2}$; and if $G=\mathrm{GU}_{n}(q)$ then $|C_{G}(g)|<2^{E}q^{d^{\prime}}\leq q^{d^{\prime}+E}$.
@@
Let now $g_{f}$ be the restriction of $g$ to $V_{f}$. Assume $f$ has degree $m$ and $V_{f}$ has dimension $k$. Note that $g_{f}$ acts $\mathbb{F}_{q^{m}}$-linearly; let $u_{f}$ be the unipotent part of $g_{f}$ as an element of $\mathrm{GL}_{k/m}(q^{m})$. Write $S=S_{f,f^{*}}$, $g^{\prime}=g_{f,f^{*}}$, $u=u_{f}$, for ease of notation. Then $C_{S}(g^{\prime})\cong C_{L}(u)$, where $L=\mathrm{GU}_{k/m}(q^{m/2})$ if $f=f^{*}$ and $\deg(f)>1$; $L=\mathrm{GL}_{k/m}(q^{m})$ if $f\neq f^{*}$; $L=\mathrm{Sp}_{k}(q)$ if $f=f^{*}$ and $\deg(f)=1$ (i.e., $f=X\pm 1$). Assume next $g$ is as in (ii). Then $g$ is semisimple, and so $u=1$ and $C_{L}(u)=L$. In particular, we see that the bound follows from Lemma 7.9, noting that the number of $f$ with $f=f^{*}$ and $f\neq X+1$ and $V_{f}\neq 0$ is at most $E_{2}$ (since each such $f$ has degree at least 22). Assume now $g$ is as in (iii), so $u$ may be nontrivial. The bound follows from (i), noting that the number of rational canonical blocks of $g^{\prime}$ is at least the number of Jordan blocks of $u$. (More precisely, if $f=f^{*}$ then it is equal to the number of Jordan blocks of $u$, and otherwise it is twice that number.)
@@
Let us then focus on Table 7.1. We denote by $\Phi_{m}(X)$ the $m$-th cyclotomic polynomial, and by $\Phi^{*}_{m}(r)$ the product of all primitive prime divisors (ppds) of $r^{m}-1$. Recall that by Zsigmondy’s theorem, $\Phi^{*}_{m}(r)>1$ unless $m=2$ and $r+1$ is a 22-power, or $(m,r)=(6,2)$. (The notation $\Phi^{*}_{m}(r)$, which is borrowed from [27], is not accidental, in the sense that $\Phi^{*}_{m}(r)$ always divides the $m$-th cyclotomic polynomials $\Phi_{m}$ evaluated at $r$.) We also denote $\Phi^{\prime}_{8}=r+\sqrt{2r}+1$, $\Phi^{\prime}_{12}=r+\sqrt{3r}+1$, $\Phi^{\prime}_{24}=r^{2}+\sqrt{2r^{3}}+r+\sqrt{2r}+1$.
@@
Note that in some cases the order of $x_{i}$ is 11. For $x_{1}$, this happens when $S$ is $\mathrm{PSL}_{6}(2)$ or $\mathrm{P}\Omega^{+}_{8}(2)$. For $x_{2}$, this happens when $S$ is $\mathrm{PSL}_{7}(2)$, $\mathrm{PSp}_{4}(r)$ (and $r+1$ is a 22-power), $\mathrm{PSp}_{8}(2)$, $\mathrm{P}\Omega^{+}_{12}(2)$, or $\mathrm{P}\Omega^{-}_{8}(2)$. In particular, in all these cases we regard $x_{i}$ as being not defined.
@@
(i) The last statement (“and so…”) follows from the first. The first statement follows from inspection of the Schur multiplier $M$ of $S$ and, in groups of Lie type, from the fact that each prime divisor of $\Phi^{*}_{c}(r)$ is congruent to 11 mod $c$. Let us handle for example $S=\mathrm{PSL}_{m}(r)$. Each prime divisor of $M$ divides $m$, and each prime divisor of $|x_{1}|$ is at least $m+1$, so $(|x_{1}|,M)=1$. As for $x_{2}$, the only possible exception occurs when $m$ is a prime divisor of $\Phi^{*}_{m-1}(r)$. But then by definition of ppd, we have that $m$ does not divide r1r-1 and so does not divide $M$.
@@
(iii) Assume first $S$ is classical, let $\tilde{S}$ be the universal covering group of $S$ and let $\tilde{x}_{i}$ be a lift of $x_{i}$ as in (i). We readily see that $\tilde{x}_{i}$ has distinct eigenvalues on the natural module (in which case $x_{i}$ is regular) except possibly in the following cases: $i=2$ and $S=\mathrm{PSU}_{m}(r),\mathrm{PSp}_{2m}(r),\mathrm{P}\Omega^{\pm}_{2m}(r),\mathrm{P}\Omega_{2m+1}(r)$; or $i=1$ and $S=\mathrm{P}\Omega^{+}_{2m}(r)$. In the case $i=2$ and $S=\mathrm{P}\Omega^{-}_{2m}(r)$, or $i=1$ and $S=\mathrm{P}\Omega^{+}_{2m}(r)$, we have that $\tilde{x}_{i}$ centralizes a 22-space and acts irreducibly on a complement, so $x_{i}$ is regular. In the case $i=2$ and $S=\mathrm{PSp}_{2m}(r)$, or $S=\mathrm{PSU}_{m}(r)$ with $m\equiv 3\pmod{4}$, $\tilde{x}_{i}$ centralizes a 22-space and so $x_{i}$ is not regular. Let us go through the remaining cases. If $S=\mathrm{PSU}_{m}(r)$ with $m$ even or $m\equiv 1\pmod{4}$, a straightforward calculation shows that $\tilde{x}_{2}$ has distinct eigenvalues if and only if $m\equiv 0,1\pmod{4}$ (that is, if and only if a totally singular irreducible $\mathbb{F}_{r}\langle\tilde{x}_{2}\rangle$-submodule has even dimension). On the other hand, if $S=\mathrm{P}\Omega^{+}_{2m}(r),\mathrm{P}\Omega_{2m+1}(r)$ then $\tilde{x}_{2}$ has distinct eigenvalues if and only if $m$ is odd. This concludes the proof if $S$ is classical.
@@
(vii) If $S$ is sporadic we use [12], and if $S$ is of Lie type this is a straightforward check. We only note that the parameter $\varepsilon$ for $S=G_{2}(r)$ ensures that $|x_{1}|$ is not divisible by 33. ∎
@@
Let us recall that if $x_{i}$ is regular, then the value $N_{L}(\langle y_{i}\rangle)/C_{L}(y_{i})$ can be read off from the Weyl group, as follows. Assume for convenience that $L=X^{F}$ is of simply connected type, so by Lemma 7.12(iv) we have $N_{L}(\langle y_{i}\rangle)/C_{L}(y_{i})=N_{X^{F}}(T^{F})/T^{F}$. Write $T=R_{w}$ where $R$ is a fixed $F$-stable maximal torus and $w\in W:=N_{X}(R)/R$ (see [63, Section 25.1] for the notation used). Since in our case $N_{X^{F}}(T^{F})=N_{X^{F}}(T)$, we have that $|N_{X^{F}}(T^{F})/T^{F}|=|C_{W}(Fw)|$, where we view FwFw as an element of the coset FWFW of $W\rtimes\langle F\rangle$; see for example [63, Proposition 25.3].